Sam has $6.00 in coins. Half of them are quarters, a third are dimes, and the rest are nickels. How many of each coin does he have?
Answer by lwsshak3(11628)

Sam has $6.00 in coins. Half of them are quarters, a third are dimes, and the rest are nickels. How many of each coin does he have?
***
let x=total number of coins
x/2=number of quarters
x/3=number of dimes
x-x/2-x/3=x-3/6-2/6=1/6=number of nickels
.25(x/2)+.10(x/3)+.05(x/6)=6.00
lcd:6
.25*3x+.10*2x+.05x=36.00
.75x+.20x+.05x=36.00
x=36
x/2=18
x/3=12
x/6=6
How many of each coin does he have?
number of quarters=18
number of dimes=12
number of nickels=6